Suppose that P(A) = 0.2 and P(B) = 0.5 and P(A ∪ B) = 0.6. Find P(A' ∪ B' )
Answer #1

By the Formula:

P(A U B) = P(A) + P(B) - P(A \cap B)

0.6 = 0.2 + 0.5 - P(A \cap B)

P(A \cap B) = 0.1

Hence,

P(A' U B')

= 1 - P(A \cap B)

= 1 - 0.1

= 0.9